# ID-Booth: Identity-consistent Face Generation with Diffusion Models

### v4

@@ I Introduction
State-of-the-art generative models are currently dominated by diffusion-based techniques, which offer unparalleled synthesis capabilities in terms of quality and diversity of the generated data, while enabling synthesis guided by text prompts [40]. Recently, diffusion models have also been utilized to produce datasets suitable for face recognition tasks, i.e., containing images of multiple identities with multiple samples each. To this end, approaches rely on identity-conditioning [34, 51, 52] and fine-tuning [42, 35] of pretrained diffusion models. Nevertheless, most solutions focus mainly on image reconstruction during training, resulting in poor consistency between the desired identities and the generated ones. To address this issue, PortraitBooth [35] recently extended the fine-tuning DreamBooth [42] method with an identity-based training objective. However, the proposed solution only considers the identity similarity of input samples and the generated samples during training. In turn, it tends to overfit on input identity features, including undesired characteristics, e.g., the pose, age, hair, accessories, thereby reducing the diversity of generated images.
@@
In this paper, we present a solution for the outlined issues, in the form of a new generative framework, called ID-Booth. The proposed framework entails three main components, including $(i)$ a denoising network that produces data based on input noise, $(ii)$ a Variational Auto-Encoder (VAE) that maps images to and from a more efficient latent space on which the denoising network operates, and $(iii)$ a text encoder that enables prompt-based conditioning of the denoising network. The proposed framework utilizes a novel triplet identity objective, which considers both positive and negative identity samples during training, to facilitate the generation of identity-consistent images while retaining the synthesis capabilities of pretrained models. Throughout the experiments, we explore the suitability of ID-Booth for addressing privacy concerns by generating diverse synthetic in-the-wild images of identities from the Tufts Face Database [33], which contains images gathered in a constrained setting with subject consent, as shown in Figure 1. We perform fine-tuning of a state-of-the-art diffusion model conditioned on diverse prompts and compare synthesis results with DreamBooth [42] and PortraitBooth [35] in terms of image quality, fidelity and diversity as well as intra-identity consistency and inter-identity separability. Furthermore, we investigate the real-world utility of the produced synthetic samples for augmenting existing datasets to train modern face recognition models in a privacy-preserving manner. We demonstrate that our fine-tuning framework enables the generation of more diverse synthetic samples with better intra-identity consistency and inter-identity separability. As showcased by improved recognition performance, across five real-world verification benchmarks, this makes our approach more suitable for augmenting small-scale training datasets than DreamBooth [42] or PortraitBooth [35]. Overall, the paper makes the following contributions:
@@ II Related work
Recently, Boutros et al. [4] achieved the generation of identity-specific images with latent diffusion models by conditioning the denoising network on face recognition features. The proposed contextual partial dropout also prevented overfitting on identities and enabled control over inter-identity separability and intra-identity diversity. Differently, more recent approaches relied on pretrained diffusion models [40] rather than training the models from scratch. Ruiz et al. [42] presented the DreamBooth method that can associate a new identity to a rare text token through fine-tuning on images of the identity. During training, face images generated by the pretrained model are also used to preserve prior synthesis capabilities. Arc2Face [34] instead replaces the identity token with recognition features and fine-tunes the model on a large-scale dataset. The textual-part of the prompt is frozen, so that control is tied primarily to the identity features, thus enabling more consistent generation of input identities. However, this comes at the cost of losing powerful prompt-based control. The recent IP-Adapter [52] has also been modified to use identity features as the condition, while retaining control of text prompts through decoupled cross-attention. InstantID [51] extends these capabilities by incorporating spatial control with an auxiliary ControlNet-based [54] module conditioned on facial landmarks and features. Despite advancements, identity consistency remained problematic, as the identity aspect was not considered in training objectives. To address this, Peng et al. [35] introduced PortraitBooth, which incorporates an identity-based objective into the fine-tuning of DreamBooth [42]. However, the solution only relies on the identity similarity of training images and generated noisy images, despite the success of more refined objectives on face recognition tasks [49]. As a result, the approach can overfit even on undesired characteristics of training identities, e.g., their pose, age or face accessories, which lowers the diversity of produced images. In contrast, our proposed ID-Booth framework utilizes a triplet objective that relies on the identity similarity between generated images and both training images (i.e., positive samples) and prior images produced by the initial model (i.e., negative samples). This enables better identity consistency, while better retaining synthesis capabilities of pretrained latent diffusion models.
@@ IV Experiments and Results
Evaluation methodology. We evaluate the suitability of the proposed ID-Booth framework by comparing its synthesis capabilities to those of DreamBooth [42] and PortraitBooth [35]. Other diffusion-based frameworks that produce identity-specific images, e.g., Arc2Face [34] and InstantID [51], are not considered as they are trained on large-scale web-scraped face recognition datasets, without the consent of subjects. Meanwhile, our experiments entail fine-tuning on a limited amount of images from TFD [33] gathered with suitable consent. To evaluate the produced images, we compare them to the diverse real-world images of FFHQ [24]. Here, we consider either entire images or only the face regions of a resolution $112\times 112$, which are aligned and cropped based on face landmarks detected by MTCNN [53]. The quality of images is then determined with the use of Fréchet Distance [17] and Kernel Distance [2], measured on features extracted with the pretrained DINOv2-ViT-L/14 model [32] rather than the typical Inception-v3 [47] model, which has been shown to be unsuitable, due to poor correlation with human evaluators [46] and the limitations of the ImageNet dataset [8]. We also evaluate the fidelity and diversity of images separately, with the use of Density and Coverage [31], measured on features of DINOv2-ViT-L/14 [32]. To compute these scores we utilize the generated datasets with $100$ samples per identity and compare them to $10.000$ samples from FFHQ [24]. In addition, we analyze the intra-identity diversity of samples with the Vendi score [12] computed on the extracted features, which differently from previous measures does not require a reference dataset. Similarly, we rely on the Certainty Ratio Face Image Quality Assessment (CR-FIQA) [3] to evaluate the quality of each face image through relative classifiability with a pretrained ResNet-101 [15]. We also analyze intra-identity diversity by evaluating the pitch, yaw and roll of faces in the images with the 6DRepNet [16] head pose estimator.
@@
Figure 3: Comparison of generated image samples. ID-Booth facilitates better identity consistency than DreamBooth [42] and better image diversity than PortraitBooth [35], which limits the variety of facial features and poses.
@@ IV-A Evaluation of generated images
Image quality. We begin our evaluation by assessing the overall quality of images produced by either the proposed ID-Booth framework or its two main competitors, DreamBooth [42] and PortraitBooth [35]. To this end, we utilize the Fréchet Distance [17] and Kernel Distance [2] computed between features extracted with DINOv2-ViT-L/14 [32] from synthetic images and features extracted from real-world images of the FFHQ [24] dataset. From results reported in Table I and samples in Figures 1 and 3, we can discern that with the SD-2.1 model and our defined prompts we can generate images that better match the quality of in-the-wild FFHQ [24] images than the real-world images of TFD [33], which were gathered in a constrained environment. Comparing results of the different fine-tuning methods, we see that our ID-Booth framework achieves the best quality results, scoring closest to the non fine-tuned model, while enabling identity-specific generation. This is the case both when evaluating entire images or only the cropped face regions. Interestingly, both approaches with identity-based training objectives (i.e., PortraitBooth [35] and our ID-Booth) also generate images of a better quality than DreamBooth [42].
@@
In addition, we evaluate the quality of each face region with CR-FIQA [3]. Here, however, high face quality is not necessarily as desired as having a mix of high and low quality images, which can lead to better performing recognition models. This difference can also be observed on real-world datasets, where in-the-wild images of FFHQ [24] achieve a lower mean but higher standard deviation than constrained images of TFD [33]. Similarly, compared to DreamBooth [42] and PortraitBooth [35], our ID-Booth framework achieves a higher standard deviation but lower mean of CR-FIQA [3] scores, that are closer to those of diverse samples of the non fine-tuned models.
@@
Image fidelity and diversity. Next, we analyze the produced images in terms of fidelity, i.e., the degree to which they resemble real samples, and diversity, i.e., how well they cover the variability of real samples [43]. To this end, we rely on Density and Coverage [31], respectively, reported in Table I, in addition to qualitative samples in Figures 3 and 4. Comparing different datasets with FFHQ [24], we can observe that images of TFD [33] lack the fidelity and diversity expected of in-the-wild images. Synthetic samples produced by the non-finetuned SD-2.1 model with diverse prompts offer a notable improvement in these areas. In comparison, fine-tuning approaches achieve a higher fidelity of entire images and face regions, but crucially also result in lower diversity of face regions. Among the approaches, DreamBooth [42] scores the highest in terms of density (i.e., fidelity), while our ID-Booth achieves the highest coverage (i.e., diversity) on both entire images and detected face regions. This can also be observed in Figures 3 and 4, where ID-Booth generates images that offer more consistent identities, while allowing for a larger variety of poses, ages, accessories and other facial features, unlike DreamBooth [42] or PortraitBooth [35].
@@
Intra-identity diversity. To further investigate the produced images, we also analyze the intra-diversity of samples with the per-class Vendi score [12]. As reported in Table I, synthetic images generated by all fine-tuning methods offer more intra-identity diversity than the constrained samples of TFD [33]. Our ID-Booth achieves the largest intra-identity diversity among the fine-tuning approaches, both of entire images and only face regions, while ensuring better identity consistency. This can be seen in Figure 4, where ID-Booth samples of the same identity contain a larger variety of poses and face accessories. To obtain deeper insight, we also analyze the pitch, yaw and roll of faces with the 6DRepNet [16] head pose estimator. In Table II we report the mean and standard deviation of standard deviation values obtained from pose distributions of each identity. Results reveal that ID-Booth generates samples with the largest variety of poses per identity, especially in terms of pitch and roll, which represents an important aspect of overall intra-identity diversity. In comparison, DreamBooth [42] and PortraitBooth [35] often default to more front-facing poses, as seen in Figures 3 and 4.
@@ IV-B Recognition-based experiments
Identity consistency and separability. To determine the suitability of generated images for augmenting recognition datasets we must also examine the consistency and separability of identities in the images. To this end, we form genuine and imposter distributions either only among synthetic identities or between synthetic and real-world identities, based on the similarity of features extracted with the pretrained ArcFace-based recognition model [9]. From verification results in Table III that describe these distributions, we can discern that both PortraitBooth [35] and ID-Booth achieve better identity consistency and separability than DreamBooth [42], due to the use of identity-based training objectives. This is the case both either among synthetic or between synthetic and real identities across all verification metrics. The only exception is the FNMR1000 score in the latter scenario, which can likely be attributed to a handful of outliers, especially when considering the lower FNMR100 scores and improved FDR values. Compared to PortraitBooth [35], our ID-Booth achieves lower FNMR scores among synthetic identities along with lower FMR and FNMR scores between synthetic and real identities, indicating fewer outliers. It also achieves a notably higher FDR score in the second scenario, which in combination with above observations signifies better intra-identity consistency and inter-identity separability. This is further supported by ID-Booth samples in Figure 4, which showcase better consistency between generated and real identities or among different synthetic samples of the same identity. Overall, these results highlight crucial characteristics of ID-Booth that demonstrate its suitability for augmenting existing datasets in a privacy-preserving manner by producing identity-consistent in-the-wild images of real-world identities from the training dataset collected with subject consent.
@@
We first explore augmentation with synthetic datasets consisting of $21$ samples per identity, denoted as SD-2.1 (21), which matches the scale of TFD [33]. Results reported in Table IV reveal that this form of augmentation enables the training of drastically better performing recognition models compared to training on only real-world samples of TFD [33]. Among the fine-tuning approaches, our ID-Booth framework achieves the highest overall accuracy, with improvements over DreamBooth [42] and PortraitBooth [35] being especially evident on the AgeDB-30 [30] and CP-LFW [55] benchmarks. In total, the use ID-Booth for augmentation results in a $6.1\%$ average accuracy increase over the recognition model trained on the non-augmented dataset. In the second set of experiments, we investigate the augmentation capabilities of larger synthetic datasets consisting of $100$ samples per identity, denoted as SD-2.1 (100). From results in Table IV we can discern notable accuracy improvements across with all fine-tuning approaches, showcasing their scalability. Similar to previous results, our ID-Booth framework achieves higher overall verification accuracy than DreamBooth [42] or PortraitBooth [35], with a total average augmentation improvement of $7.9\%$. As before, the most noticeable improvement is observed on cross-age benchmarks (i.e., AgeDB-30 [30] and CA-LFW [56]). Overall, the accuracy improvements obtained with our ID-Booth framework can be attributed to the novel triplet identity objective, which facilitates better identity consistency and higher intra-identity diversity, especially in terms of age and pose, than either DreamBooth [42] or PortraitBooth [35].

### v5

@@ I Introduction
State-of-the-art generative models are currently dominated by diffusion-based techniques, which offer unparalleled synthesis capabilities in terms of quality and diversity of the generated data, while enabling synthesis guided by text prompts [40]. Recently, diffusion models have also been utilized to produce datasets suitable for face recognition tasks, i.e., containing images of multiple identities with multiple samples each. To this end, approaches rely on identity-conditioning [34, 51, 52] and fine-tuning [42] of pretrained diffusion models. Nevertheless, most solutions focus mainly on image reconstruction during training, resulting in poor consistency between the desired and the generated identities. To address this issue, PortraitBooth [35] proposed an additional identity-based training objective, which can be used to extend the fine-tuning DreamBooth [42] method. However, the solution only considers the identity similarity of input samples and the generated samples during training. In turn, it tends to overfit on input identity features, including undesired characteristics, e.g., the pose, age, hair, accessories, thereby reducing the diversity of generated images.
@@
In this paper, we present a solution for the outlined issues, in the form of a new generative framework, called ID-Booth. The proposed framework entails three main components, including $(i)$ a denoising network that produces data based on input noise, $(ii)$ a Variational Auto-Encoder (VAE) that maps images to and from a more efficient latent space on which the denoising network operates, and $(iii)$ a text encoder that enables prompt-based conditioning of the denoising network. The proposed framework utilizes a novel triplet identity objective, which considers both positive and negative identity samples during training, to facilitate the generation of identity-consistent images while retaining the synthesis capabilities of pretrained models. Throughout the experiments, we explore the suitability of ID-Booth for addressing privacy concerns by generating diverse synthetic in-the-wild images of identities from the Tufts Face Database [33], which contains images gathered in a constrained setting with subject consent, as shown in Figure 1. We perform fine-tuning of a state-of-the-art diffusion model conditioned on diverse prompts and compare synthesis results with DreamBooth [42] and a PortraitBooth-based [35] version of it in terms of image quality, fidelity and diversity as well as intra-identity consistency and inter-identity separability. Furthermore, we investigate the real-world utility of the produced synthetic samples for augmenting existing datasets to train modern face recognition models in a privacy-preserving manner. We demonstrate that our fine-tuning framework enables the generation of more diverse synthetic samples with better intra-identity consistency and inter-identity separability. As showcased by improved recognition performance, across five real-world verification benchmarks, this makes our approach more suitable for augmenting small-scale training datasets than existing solutions [42, 35]. Overall, the paper makes the following contributions:
@@ II Related work
Recently, Boutros et al. [4] achieved the generation of identity-specific images with latent diffusion models by conditioning the denoising network on face recognition features. The proposed contextual partial dropout also prevented overfitting on identities and enabled control over inter-identity separability and intra-identity diversity. Differently, more recent approaches relied on pretrained diffusion models [40] rather than training the models from scratch. Ruiz et al. [42] presented the DreamBooth method that can associate a new identity to a rare text token through fine-tuning on images of the identity. During training, face images generated by the pretrained model are also used to preserve prior synthesis capabilities. Arc2Face [34] instead replaces the identity token with recognition features and fine-tunes the model on a large-scale dataset. The textual-part of the prompt is frozen, so that control is tied primarily to the identity features, thus enabling more consistent generation of input identities. However, this comes at the cost of losing powerful prompt-based control. The recent IP-Adapter [52] has also been modified to use identity features as the condition, while retaining control of text prompts through decoupled cross-attention. InstantID [51] extends these capabilities by incorporating spatial control with an auxiliary ControlNet-based [54] module conditioned on facial landmarks and features. Despite advancements, identity consistency remained problematic, as the identity aspect was not considered in training objectives. To address this, Peng et al. [35] introduced PortraitBooth, a method that incorporates an identity-based objective into the training process, which can also be applied to the fine-tuning of DreamBooth [42]. However, the solution only relies on the identity similarity of training images and generated noisy images, despite the success of more refined objectives on face recognition tasks [49]. As a result, the approach can overfit even on undesired characteristics of training identities, e.g., their pose, age or face accessories, which lowers the diversity of produced images. In contrast, our proposed ID-Booth framework utilizes a triplet objective that relies on the identity similarity between generated images and both training images (i.e., positive samples) and prior images produced by the initial model (i.e., negative samples). This enables better identity consistency, while better retaining synthesis capabilities of pretrained latent diffusion models.
@@ IV Experiments and Results
Evaluation methodology. We evaluate the suitability of the proposed ID-Booth framework by comparing its synthesis capabilities to those of DreamBooth [42] and a version of it extended with the PortraitBooth [35] identity objective. Other diffusion-based frameworks that produce identity-specific images, e.g., Arc2Face [34] and InstantID [51], are not considered as they are trained on large-scale web-scraped face recognition datasets, without the consent of subjects. Meanwhile, our experiments entail fine-tuning on a limited amount of images from TFD [33] gathered with suitable consent. To evaluate the produced images, we compare them to the diverse real-world images of FFHQ [24]. Here, we consider either entire images or only the face regions of a resolution $112\times 112$, aligned and cropped based on face landmarks detected by MTCNN [53]. The quality of images is then determined with Fréchet Distance [17] and Kernel Distance [2], measured on features extracted with the pretrained DINOv2-ViT-L/14 model [32] rather than the typical Inception-v3 [47] model, which has been shown to be unsuitable, due to poor correlation with human evaluators [46] and the limitations of the ImageNet dataset [8]. We also evaluate the fidelity and diversity of images separately, with the use of Density and Coverage [31], measured on features of DINOv2-ViT-L/14 [32]. To compute these scores we utilize the generated datasets with $100$ samples per identity and compare them to $10.000$ samples from FFHQ [24]. In addition, we analyze the intra-identity diversity of samples via the extracted features with the Vendi score [12], which differently from previous measures does not require a reference dataset. Similarly, we rely on the Certainty Ratio Face Image Quality Assessment (CR-FIQA) [3] to evaluate the quality of each face image through relative classifiability with a pretrained ResNet-101 [15]. We also analyze intra-identity diversity by evaluating the pitch, yaw and roll of faces in the images with the 6DRepNet [16] head pose estimator.
@@
Figure 3: Comparison of generated image samples. ID-Booth facilitates better identity consistency than DreamBooth [42] and better image diversity than when utilizing the PortraitBooth [35] identity objective, which can limit the variety of facial features and poses.
@@ IV-A Evaluation of generated images
Image quality. We begin our evaluation by assessing the overall quality of images produced by either the proposed ID-Booth framework or its two competitors, the base DreamBooth [42] and a version of it extended with the PortraitBooth [35] identity objective (denoted as PortraitBooth for brevity). To this end, we utilize the Fréchet Distance [17] and Kernel Distance [2] computed between features extracted with DINOv2-ViT-L/14 [32] from synthetic images and features extracted from real-world images of the FFHQ [24] dataset. From results reported in Table I and samples in Figures 1 and 3, we can discern that with the SD-2.1 model and our defined prompts we can generate images that better match the quality of in-the-wild FFHQ [24] images than the real-world images of TFD [33], which were gathered in a constrained environment. Comparing results of the different fine-tuning methods, we see that our ID-Booth framework achieves the best quality results, scoring closest to the non fine-tuned model, while enabling identity-specific generation. This is the case both when evaluating entire images or only the cropped face regions. Interestingly, both identity-based training objectives of either PortraitBooth [35] or ID-Booth improve the image quality of the base DreamBooth [42].
@@
In addition, we evaluate the quality of each face region with CR-FIQA [3]. Here, however, high face quality is not necessarily as desired as having a mix of high and low quality images, which can lead to better performing recognition models. This difference can also be observed on real-world datasets, where in-the-wild images of FFHQ [24] achieve a lower mean but higher standard deviation than constrained images of TFD [33]. Similarly, compared to DreamBooth [42] and the PortraitBooth-based [35] version, our ID-Booth framework achieves a higher standard deviation but lower mean of CR-FIQA [3] scores, that are closer to those of diverse samples of the non fine-tuned models.
@@
Image fidelity and diversity. Next, we analyze the produced images in terms of fidelity, i.e., the degree to which they resemble real samples, and diversity, i.e., how well they cover the variability of real samples [43]. To this end, we rely on Density and Coverage [31], respectively, reported in Table I, in addition to qualitative samples in Figures 3 and 4. Comparing different datasets with FFHQ [24], we can observe that images of TFD [33] lack the fidelity and diversity expected of in-the-wild images. Synthetic samples produced by the non-finetuned SD-2.1 model with diverse prompts offer a notable improvement in these areas. In comparison, fine-tuning approaches achieve a higher fidelity of entire images and face regions, but crucially result in lower diversity of face regions. Among the approaches, DreamBooth [42] scores the highest in terms of density (i.e., fidelity), while ID-Booth achieves the highest coverage (i.e., diversity) on both entire images and detected face regions. This can also be observed in Figures 3 and 4, where ID-Booth offers more consistent identities, while enabling a larger variety of poses, ages, accessories and other facial features than DreamBooth [42] or the PortraitBooth-based [35] version of it.
@@
Intra-identity diversity. To further investigate the produced images, we also analyze the intra-diversity of samples with the per-class Vendi score [12]. As reported in Table I, synthetic images generated by all fine-tuning methods offer more intra-identity diversity than the constrained samples of TFD [33]. Our ID-Booth achieves the largest intra-identity diversity among the fine-tuning approaches, both of entire images and only face regions, while ensuring better identity consistency. This can be seen in Figure 4, where ID-Booth samples of the same identity contain a larger variety of poses and face accessories. To obtain deeper insight, we also analyze the pitch, yaw and roll of faces with the 6DRepNet [16] head pose estimator. In Table II we report the mean and standard deviation of standard deviation values obtained from pose distributions of each identity. Results reveal that ID-Booth generates samples with the largest variety of poses per identity, especially in terms of pitch and roll, which represents an important aspect of overall intra-identity diversity. In comparison, DreamBooth [42] and the PortraitBooth-based [35] version often default to more front-facing poses, as seen in Figures 3 and 4.
@@ IV-B Recognition-based experiments
Identity consistency and separability. To determine the suitability of generated images for augmenting recognition datasets we must also examine the consistency and separability of identities in the images. To this end, we form genuine and imposter distributions either only among synthetic identities or between synthetic and real-world identities, based on the similarity of features extracted with the pretrained ArcFace-based recognition model [9]. From verification scores in Table III that describe these distributions, we can discern that the use of identity-based training objectives from either ID-Booth or PortraitBooth [35] results in better identity consistency and separability than the base DreamBooth [42]. This is the case both either among synthetic or between synthetic and real identities across all verification measures. The only exception is the FNMR1000 score in the latter scenario, which can likely be attributed to a handful of outliers, especially when considering the lower FNMR100 scores and improved FDR values. Compared to the PortraitBooth-based [35] approach, ID-Booth achieves lower FNMR scores among synthetic identities along with lower FMR and FNMR scores between synthetic and real identities, indicating fewer outliers. In combination with a higher FDR score in the second scenario, this signifies better intra-identity consistency and inter-identity separability. ID-Booth samples in Figure 4 further support these observations with better consistency between generated and real identities or among different samples of the same identity. Overall, these results highlight crucial characteristics of ID-Booth, demonstrating its suitability for augmenting existing datasets in a privacy-preserving manner by producing identity-consistent in-the-wild images of real-world identities from the training dataset collected with subject consent.
@@
We first explore augmentation with synthetic datasets consisting of $21$ samples per identity, denoted as SD-2.1 (21), which matches the scale of TFD [33]. Results reported in Table IV reveal that this form of augmentation enables the training of drastically better performing recognition models compared to training on only real-world samples of TFD [33]. Among the fine-tuning approaches, our ID-Booth framework achieves the highest overall accuracy, with improvements being especially evident on the AgeDB-30 [30] and CP-LFW [55] benchmarks. In total, using ID-Booth for augmentation facilitates a $6.1\%$ average accuracy increase over the recognition model trained on the non-augmented dataset. In the second set of experiments, we investigate the augmentation capabilities of larger synthetic datasets consisting of $100$ samples per identity, denoted as SD-2.1 (100). From results in Table IV we can discern notable accuracy improvements across with all fine-tuning approaches, showcasing their scalability. Similar to previous results, our ID-Booth framework achieves higher overall verification accuracy than DreamBooth [42] or the PortraitBooth-based [35] version, with a total average augmentation improvement of $7.9\%$. As before, the most noticeable improvement is observed on cross-age benchmarks (i.e., AgeDB-30 [30] and CA-LFW [56]). Overall, the accuracy improvements obtained with our ID-Booth framework can be attributed to the novel triplet identity objective, which enables better identity consistency and higher intra-identity diversity, especially in terms of age and pose, than either DreamBooth [42] or the PortraitBooth-based [35] version.